Sixty percent of the customers of a fast food chain order a hamburger, french fries and a drink. If a random sample of 15 cash register receipts is selected, what is the probability that 10 or more will show that the above three food items were ordered?
Answer by edjones(7569)

Let a=.6 and b=.4
Take the 1st six terms of (a+b)^15 which are
find the coefficients by the combination formula nCr on your calculator (easiest).
a^15=.0005
15a^14b=.0047
105a^13b^2=.0219
455a^12b^3=.0634
1365a^11b^4=.1268
3003a^10b^5=.1859
Sum of the above numbers .4032 answers.
.
Ed